If an automobile gets an average of 27 miles per gallon on a trip and the standard deviation is 3 miles per gallon, find the probability that on a randomly selected trip, the automobile will get between 21 and 30 miles per gallon. Assume the variable is normally distributed.


1
Expert's answer
2021-03-31T15:29:42-0400

Let "X=" distance per gallon on a trip in miles: "X\sim N(\mu, \sigma^2)"

Given "\mu=27" miles per gallon, "\sigma=3" miles per gallon.

Then


"P(21<X<30)=P(X<30)-P(X\leq 21)"

"=P(Z<\dfrac{30-27}{3})-P(Z\leq \dfrac{21-27}{3})"

"=P(Z<1)-P(Z\leq -2)"

"\approx0.841345-0.022750\approx0.8186"

The probability that on a randomly selected trip, the automobile will get between 21 and 30 miles per gallon is 0.8186.